Fringes found with Hb and Yg on source 1921-293 (Dave)
NB: module is only 1500 GB, whereas expected size of experiment is 1580 GB. This was not noticed until data was on the module. Last 80 GB will need to be transported another way. (Dave) - Liza says the solution will be to swap to the other bank, since no one is a the station today (Lucia).
2015.167.11:43:34 ALARM: Delay through DBBC has changed. Monitor for stability and reset DBBC if it drifts.
11:56 UT - dbbc rebooted and recording resumed, missed scans 167-1143 through 167-1153 (Ross).